# Hollow toroidal rotation profiles in strongly electron heated H-mode plasmas in the ASDEX Upgrade tokamak

### 05 Mar 2026 v1

@@ 1 Introduction
In the modeling, following an approach discussed in a very similar manner in Ref. [Zimmermann_NF_Isotope], an experimental boundary condition is imposed at a normalized toroidal flux radius of approximately $\rho_{\varphi}=0.8$. Variations in the precise radial location of this boundary condition have only a minor effect, provided that the analyzed profiles remain sufficiently smooth and that the boundary condition avoids regions affected by edge-localized modes (ELMs). The flux-surface-averaged radial transport of toroidal angular momentum is expressed as
@@
$\Gamma_{\varphi}=-mnR\left(\chi_{\varphi}\,\frac{\partial v_{\varphi}}{\partial r}-V_{\mathrm{c}}\,v_{\varphi}\right)+\Pi_{\text{int}}\,$, (1)
@@
where $m$ denotes the mass of the main ion species, $n$ its density, and $R$ the flux-surface-averaged local major radius. The coefficient $\chi_{\varphi}$ represents the momentum diffusivity. The toroidal velocity is defined as $v_{\varphi}=\langle R\Omega_{\varphi}\rangle$, where $\langle\cdot\rangle$ denotes a flux-surface average and $\Omega_{\varphi}$ is the toroidal component of the angular rotation frequency. For sufficiently weak poloidal plasma rotation, as is typically the case in the plasma core of the discharges considered here, $\Omega_{\varphi}$ can be assumed to be constant on a given flux surface [lebschy2017measurement]. If poloidal rotation becomes non-negligible, the experimentally measured toroidal rotation must be corrected prior to comparison with the parallel velocity used in the modeling framework. The operator $\partial/\partial r$ denotes differentiation with respect to the minor radius $r$, $V_{\mathrm{c}}$ is the convective velocity, and $\Pi_{\text{int}}$ represents the residual stress contribution. This residual stress gives rise to an intrinsic torque defined as $\tau_{\text{int}}=-\partial V/\partial r\,\Pi_{\text{int}}$, where $V$ is the plasma volume enclosed by a given flux surface.
@@
Within the transport modeling framework, the momentum diffusivity $\chi_{\varphi}$ is linked to the experimentally inferred ion heat diffusivity $\chi_{i}$ through a linear dependence on the normalized toroidal flux coordinate $\rho_{\varphi}$. Their ratio defines the Prandtl number, $Pr=\chi_{\varphi}/\chi_{i}$, which is typically found to be of order unity. This observation supports the assumption of strong coupling between turbulent momentum and heat transport [Peeters2005, waltz2007coupled, weiland2009symmetry, peeters2006toroidal, Holod_2008, Holod_2010, Strintzi2008, Peeters2011]. The ion heat diffusivity $\chi_{i}=-Q_{i}/(n\nabla T_{i})$ (with the ion heat flux $Q_{i}$ and the ion temperature gradient $\nabla T_{i}$) is obtained from power balance calculations based on experimental measurements, ensuring that variations in turbulence intensity, such as those induced by modulation of NBI heating, are consistently included in the momentum transport modeling. The dimensionless quantity $RV_{\mathrm{c}}/\chi_{\varphi}$ is commonly referred to as the pinch number. The radial profiles of the pinch number and residual stress are prescribed using quartic polynomials and are chosen to vanish at $\rho_{\varphi}=0$ for continuity. Both convective and residual stress fluxes, are chosen to scale with the inferred $\chi_{\varphi}$ to account properly for possible modulations in the turbulence amplitude, as discussed in detail in Refs. [Zimmermann_NF_Letter, Zimmermann_NF_Isotope]. The externally applied torque and heat fluxes from NBI are calculated using the Monte Carlo code NUBEAM [pankin2004], which is embedded within the TRANSP framework [TRANSP_Reference_Paper, hawryluk1981empirical] and incorporated self-consistently into the momentum balance equation. For modeling the heat fluxes of the ECRH, the TORBEAM code is included in the TRANSP workflow [Poli2018].
@@ 2 Experimental Description
In this section, the AUG plasma discharge #29216 is introduced. It is a standard type-I ELMy H-mode plasma in a toroidal field of $B_{\varphi}=-2.5$ T, a plasma current of $I_{p}=0.6$ MA, a line averaged core electron density of $7.2\cdot 10^{19}$ m-3, with a lower-single-null, favorable-drift configuration.
@@
Figure 1 shows time traces of the most relevant physics quantities. As illustrated by the blue NBI power trace in Fig. 1(a), a single full-power steady beam ($2.4$ MW, $59$ keV) is used for plasma heating and serves as the diagnostic beam for CXRS [Viezzer2012, McDermott2017]. In addition, a reduced-power beam at approximately $1.3$ MW ($65$ keV) is modulated at a frequency of $2$ Hz for the momentum transport analysis. The discharge consists of a phase with low ECRH power to avoid impurity accumulation [Angioni_2017, Stober_2003], $P_{\mathrm{ECRH}}=0.6$ MW (from $2.0$ to $4.5$ s, hereafter referred to as the “NBI phase”). It is followed by a phase in which strong ECRH, $P_{\mathrm{ECRH}}=3.4$ MW, is applied (from $5.5$ to $7.0$ s, referred to as the “NBI+ECRH phase”). The time intervals selected for analysis are indicated by dotted and dashed vertical lines, respectively.
@@
The sawtooth (ST) inversion radius increases from $\rho_{\varphi}=0.18$ in the NBI phase to $\rho_{\varphi}=0.22$ in the NBI+ECRH phase, confirmed by analysis of electron cyclotron emission radiometry (ECE) data. This change is accounted for in the analysis later on by a slight adjustment of the analysis window and is attributed to the modified core heating and possible alterations of the current profile induced by ECRH. The sawtooth frequency exceeds $50$ Hz in both phases. Aside from sawtooth activity, no strong core MHD modes are identified in the analysis of the magnetic probe signals. ELMs occur at high frequency and, similar to sawteeth, are not temporally resolved by the CXRS diagnostics, which operate at a time resolution of $10$ ms. Consequently, the transport analysis excludes the corresponding radial regions in order to focus on turbulent transport effects. Owing to the low NBI modulation frequency of $2$ Hz, neither sawteeth nor ELMs are expected to introduce spurious higher harmonics into the modulation response.
@@
The electron density is found to increase significantly during the NBI+ECRH phase, indicating that the additional ECRH drives the plasma into an ITG-TEM mixed-mode turbulence regime characterized by an inward particle pinch [Fable_2010, Angioni_2011_gyro, Fable_2019]. This observation suggests that the applied level of ECRH power, in combination with the still substantial ion heating provided by NBI and the relatively high plasma density, is insufficient to establish a deeply TEM-dominated regime during the second discharge phase.
@@ 3 Momentum Transport Modeling
A comparison of the amplitude profiles reveals systematically lower values during the NBI+ECRH phase, suggesting either enhanced diffusive transport or reduced convective contributions. This interpretation is further supported by the comparatively flatter phase profile observed in the NBI+ECRH case, indicative of increased diffusion in the NBI+ECRH phase [Zimmermann2022]. The overall phase shift of the NBI-induced perturbation propagating through the plasma core is rather small compared to the associated error bars or to local ”dips” of the profile. This behavior is a direct consequence of the relatively low modulation frequency of $2$ Hz. As discussed later, this results in comparatively large uncertainties in the inferred Prandtl numbers, which subsequently propagate via the coupling of convective and residual stress terms to $\chi_{\varphi}$ into all extracted momentum transport coefficients.
@@
Figure 7 shows the inferred transport coefficients, with solid blue lines corresponding to the dominant NBI phase and dashed orange lines to the NBI+ECRH phase. Panel (a) presents the Prandtl number profiles, which are very similar in both phases. This behavior is consistent with previous studies, which have shown that, in the absence of strong TEM-dominated turbulence, the Prandtl number depends only weakly on plasma parameters [Strintzi2008, Kluy_2009, Zimmermann_2024]. In both cases, the Prandtl number increases with radius, in agreement with earlier work [Zimmermann_NF_Letter, Zimmermann_NF_Isotope, Zimmermann_2024]. The asymmetric error bars arise from the very flat phase profiles: the modeled phase profiles can be flattened substantially before a meaningful increase in the fitting cost function occurs, whereas any additional steepening (corresponding to lower diffusion) quickly degrades the fit quality. If highly localized variations of the Prandtl number were present, they would likely be smoothed out by the linear radial model employed here. The use of higher-order polynomial representations is deferred to future work.
@@
As noted above, the relatively large uncertainties propagate through the analysis, resulting in sizeable error bars for the inferred pinch number profiles, shown in Panel (c). In the NBI phase, the pinch number is largest in the inner core and decreases beyond mid-radius. This behavior closely resembles the profile of the logarithmic density gradient shown in Fig. 4(a) and is consistent with the theoretical concept of the Coriolis momentum pinch, which is found in theoretical [Peeters2007_PRL] and experimental studies [Tala_2009, Tala_2011, Zimmermann_2024] to scale most strongly with the logarithmic density gradient. A similar trend is observed in the NBI+ECRH phase, which exhibits slightly higher $R/L_{n_{e}}$ values in the outer core, reflected in a correspondingly larger pinch number. While these trends are in line with expectations, the relatively large error bars warrant caution in the quantitative interpretation.
@@
The observed modification of the intrinsic torque during the NBI+ECRH phase is consistent with the theoretical concept of residual stress generation arising from profile shearing effects and turbulent eddy tilting during the transition from ITG- to TEM-dominated turbulence [Camenen2011]. In particular, the substantial variations in $R/L_{n_{e}}$, see Fig. 4(a), and in $R/L_{T_{e}}$, see Panel (e), have been identified in previous studies [Angioni_PRL_2011, Grierson_PRL_2017, Hornsby2018, Zimmermann_2024] as key drivers of counter-current intrinsic torque generation. It should be noted that, from a theoretical standpoint [Camenen2011, Hornsby2018], intrinsic torque generation is expected to correlate with second-order derivatives of the kinetic profiles, which are, however, difficult to measure accurately in the experiment. Therefore, first-order derivatives are usually used as a proxy in such an analysis. A definitive interpretation of these observations would require non-linear, global gyrokinetic simulations [Hornsby2018, Wang_PRL_2009, Wang_PoP_2010, Grierson_PRL_2017], which are beyond the scope of the present work.
@@
To gain further insight into the underlying turbulence regime, the flux-tube version of the gyrokinetic GKW code [Peeters2004, Peeters2009] was employed using experimental input profiles and a time-integration scheme. This approach allows the identification of the micro-instability with the highest linear growth rate $\gamma$, its dominant mode frequency $\omega_{r}$, and enables an analysis of the mode spectra. These calculations were done at mid-radius for both discharge phases, where the impact of the counter-current intrinsic torque appears to be strongest. The resulting mode spectra are shown in Fig. 8 for the NBI phase in Panel (a) and for the NBI+ECRH phase in Panel (b). In the NBI phase, the mode frequency, see orange points, remains positive across the entire spectrum, indicating turbulence dominated by ITG modes. The growth rates, see blue crosses, are non-zero for smaller $k_{y}\,\rho_{i}$ and vanish for smaller wave lengths. In contrast, during the NBI+ECRH phase, the spectrum transitions from ITG to TEM (see the sign change in $\omega_{r}$), suggesting the presence of an ITG-TEM mixed-mode regime. It diverges as $\gamma\sim k_{y}^{2}$, which is a common feature of such linear calculations in this spectral domain with higher $k_{y}\,\rho_{s}$ contributing less to the transport. Together with the experimental observation of an enhanced inward particle pinch during the NBI+ECRH phase, this provides strong support for the assumption of mixed-mode turbulence. Further gyrokinetic or gyrofluid modeling is left for future work.
Figure 9: Cost function values of the modeling optimization for the NBI+ECRH phase (#29216, 5.57.05.5-7.0 s). The tested pinch number (at mid-radius) is shown on the x-axis and the scaling parameter of the residual stress flux on the y-axis (at mid-radius). The final model solution is indicated by a black marker. The “valley” of acceptable solutions extends predominantly toward positive pinch numbers, suggesting inward convective momentum transport.
@@
Figure 10: Comparison of the momentum fluxes for both discharge phases for #29216 for the discharge phase with dominant NBI heating (2.04.52.0-4.5 s, l.h.s.) and for the phase with NBI+ECRH (5.57.05.5-7.0 s, r.h.s.). The black markers indicate the sum of all momentum flux contributions. Values close to zero demonstrate the temporal stability and self-consistency of the modeled solutions.

### 01 Apr 2026 v2

@@ 1 Introduction
In the modeling, following an approach discussed in a very similar manner in Ref. [Zimmermann_NF_Isotope], an experimental boundary condition is imposed at a normalized toroidal flux radius of approximately $\rho_{\varphi}=0.8$. Variations in the precise radial location of this boundary condition have only a minor effect, provided that the analyzed profiles remain sufficiently smooth and that the boundary condition avoids regions affected by edge-localized modes (ELMs). Numerical experiments performed for the scenario discussed in this work, in which the boundary conditions were varied within the experimental uncertainties, show that the resulting effects are limited to the region close to the boundary and remain within the uncertainties of the inferred transport coefficients.
The flux-surface-averaged radial transport of toroidal angular momentum is expressed as
@@
$\Gamma_{\varphi}=-mnR\left(\chi_{\varphi}\,\frac{\partial v_{\varphi}}{\partial r}-V_{\mathrm{c}}\,v_{\varphi}\right)+\Pi_{\text{Rs}}\,$, (1)
@@
where $m$ denotes the mass of the main ion species, $n$ its density, and $R$ the flux-surface-averaged local major radius. The coefficient $\chi_{\varphi}$ represents the momentum diffusivity. The toroidal velocity is defined as $v_{\varphi}=\langle R\Omega_{\varphi}\rangle$, where $\langle\cdot\rangle$ denotes a flux-surface average and $\Omega_{\varphi}$ is the toroidal component of the angular rotation frequency. For sufficiently weak poloidal plasma rotation, as is typically the case in the plasma core of the discharges considered here, $\Omega_{\varphi}$ can be assumed to be constant on a given flux surface [lebschy2017measurement]. If poloidal rotation becomes non-negligible, the experimentally measured toroidal rotation must be corrected prior to comparison with the parallel velocity used in the modeling framework. The operator $\partial/\partial r$ denotes differentiation with respect to the minor radius $r$, $V_{\mathrm{c}}$ is the convective velocity, and $\Pi_{\text{Rs}}$ represents the residual stress contribution. This residual stress gives rise to an intrinsic torque defined as
$\tau_{\text{int}}=-\partial V/\partial r\,\Pi_{\text{Rs}}\text{,}$ (2)
where $V$ is the plasma volume enclosed by a given flux surface.
@@
Within the transport modeling framework, the momentum diffusivity $\chi_{\varphi}$ is linked to the experimentally inferred ion heat diffusivity $\chi_{i}$ through a linear dependence on the normalized toroidal flux coordinate $\rho_{\varphi}$. Their ratio defines the Prandtl number, $Pr=\chi_{\varphi}/\chi_{i}$, which is typically found to be of order unity. This observation supports the assumption of strong coupling between turbulent momentum and heat transport [Peeters2005, waltz2007coupled, weiland2009symmetry, peeters2006toroidal, Holod_2008, Holod_2010, Strintzi2008, Peeters2011]. The ion heat diffusivity $\chi_{i}=-Q_{i}/(n\nabla T_{i})$ (with the ion heat flux $Q_{i}$ and the ion temperature gradient $\nabla T_{i}$) is obtained from power balance calculations based on experimental measurements, ensuring that variations in turbulence intensity, such as those induced by modulation of NBI heating, are consistently included in the momentum transport modeling. The dimensionless quantity $RV_{\mathrm{c}}/\chi_{\varphi}$ is commonly referred to as the pinch number. The radial profiles of the pinch number and residual stress are prescribed using quartic polynomials and are chosen to vanish at $\rho_{\varphi}=0$ for continuity. Both convective and residual stress fluxes are chosen to scale with the inferred $\chi_{\varphi}$ to account properly for possible modulations in the turbulence amplitude, as discussed in detail in Refs. [Zimmermann_NF_Letter, Zimmermann_NF_Isotope]. The externally applied torque and heat fluxes from NBI are calculated using the Monte Carlo code NUBEAM [pankin2004], which is embedded within the TRANSP framework [TRANSP_Reference_Paper, hawryluk1981empirical] and incorporated self-consistently into the momentum balance equation. For modeling the heat fluxes of the ECRH, the TORBEAM code is included in the TRANSP workflow [Poli2018].
@@ 2 Experimental Description
In this section, the AUG plasma discharge #29216 is introduced. It is a standard type-I ELMy H-mode plasma in a toroidal field of $B_{\varphi}=-2.5$ T, a plasma current of $I_{p}=0.6$ MA, a line-averaged core electron density of $7.2\cdot 10^{19}$ m-3, with a lower-single-null, favorable-drift configuration.
@@
Figure 1 shows time traces of the most relevant physics quantities. As illustrated by the blue NBI power trace in Fig. 1(a), a single full-power steady beam ($2.4$ MW, $59$ kV) is used for plasma heating and serves as the diagnostic beam for CXRS [Viezzer2012, McDermott2017]. In addition, a reduced-power beam at approximately $1.3$ MW ($65$ kV) is modulated at a frequency of $2$ Hz for the momentum transport analysis. The discharge consists of a phase with low ECRH power to avoid impurity accumulation [Angioni_2017, Stober_2003], $P_{\mathrm{ECRH}}=0.6$ MW (from $2.0$ to $4.5$ s, hereafter referred to as the “NBI phase”). It is followed by a phase in which strong ECRH, $P_{\mathrm{ECRH}}=3.4$ MW, is applied (from $5.5$ to $7.0$ s, referred to as the “NBI+ECRH phase”). The time intervals selected for analysis are indicated by dotted and dashed vertical lines, respectively. They encompass five modulation cycles for the NBI phase and three modulation cycles for the NBI+ECRH phase. The low ensemble does not pose an issue for the analysis due to the stable background quantities, as discussed in the following.
@@
The sawtooth (ST) inversion radius increases from $\rho_{\varphi}=0.18$ in the NBI phase to $\rho_{\varphi}=0.22$ in the NBI+ECRH phase, confirmed by manual analysis of electron cyclotron emission radiometry (ECE) data. This change is accounted for later in the analysis by a slight adjustment of the analysis window, avoiding the regions found to be affected by the ST, and is attributed to the modified core heating and possible alterations of the current profile induced by ECRH. The sawtooth frequency exceeds $50$ Hz in both phases, and if they were to influence the core profiles, the analysis would average over their effects. Aside from sawtooth activity, no strong core MHD modes are identified in the analysis of the magnetic probe signals. ELMs occur at high frequency and, similar to sawteeth, are not temporally resolved by the CXRS diagnostics, which operate at a time resolution of $10$ ms. Consequently, the transport analysis excludes the corresponding radial regions in order to focus on turbulent transport effects. Owing to the low NBI modulation frequency of $2$ Hz, neither sawteeth nor ELMs are expected to introduce spurious higher harmonics into the modulation response.
@@
The electron density is found to increase significantly during the NBI+ECRH phase, indicating that the additional ECRH drives the plasma into an ITG-TEM mixed-mode turbulence regime characterized by an inward particle pinch, resulting from modifications to the microturbulence [Angioni_2005, Fable_2008, Fable_2010, angioni2011gyrokinetic, Angioni_PRL_2011]. This observation suggests that the applied level of ECRH power, in combination with the still substantial ion heating provided by NBI and the relatively high plasma density, is insufficient to establish a deeply TEM-dominated regime during the second discharge phase.
@@ 3 Momentum Transport Modeling
A comparison of the amplitude profiles reveals systematically lower values during the NBI+ECRH phase, suggesting either enhanced diffusive transport or reduced convective contributions. This interpretation is further supported by the comparatively flatter phase profile observed in the NBI+ECRH case, indicative of increased diffusion in the NBI+ECRH phase [Zimmermann2022]. The phase shift of the NBI-induced perturbation across the plasma core is small compared to both the error bars and local variations in the profile. This is a direct consequence of the relatively low modulation frequency of $2$ Hz and could be improved in future experiments by choosing higher modulation frequencies. The small phase shift observed here provides weaker constraints on the fitting. In contrast, previous work in Ref. [Zimmermann_NF_Isotope], which used a $5$ Hz modulation, achieved steeper phase profiles, better constraining the modeling of $Pr$ and correspondingly yielding smaller uncertainties in the fitted transport quantities. As discussed later, this results in comparatively large uncertainties in the inferred Prandtl numbers, which subsequently propagate via the coupling of convective and residual stress terms to $\chi_{\varphi}$ into all extracted momentum transport coefficients.
@@
Figure 7 shows the inferred transport coefficients, with solid blue lines corresponding to the dominant NBI phase and dashed orange lines to the NBI+ECRH phase. The shown error bars correspond to a statistical $1\sigma$ confidence interval around the best-fit solution, reflecting the spread of the inferred quantities due to measurement and modeling uncertainties. A more detailed discussion of the underlying methodology is provided in Ref. [Zimmermann_NF_Isotope].
Panel (a) presents the Prandtl number profiles, which are very similar in both phases. This behavior is consistent with previous studies, which have shown that, in the absence of strong TEM-dominated turbulence, the Prandtl number depends only weakly on plasma parameters [Strintzi2008, Kluy_2009, Zimmermann_2024]. In both cases, the Prandtl number increases with radius, in agreement with earlier work [Zimmermann_NF_Letter, Zimmermann_NF_Isotope, Zimmermann_2024], which suggests that the most significant scaling of the Prandtl number is with the inverse aspect ratio as a proxy for the trapped particle fraction, as discussed in more detail in Ref. [Zimmermann_2024]. The asymmetric error bars arise from the very flat phase profiles: the modeled phase profiles can be flattened substantially before a meaningful increase in the fitting cost function occurs, whereas any additional steepening (corresponding to lower diffusion) quickly degrades the fit quality. If highly localized variations of the Prandtl number were present, they would likely be smoothed out by the linear radial model employed here. The use of higher-order polynomial representations is deferred to future work.
@@
As noted above, the relatively large uncertainties propagate through the analysis, resulting in sizeable error bars for the inferred pinch number profiles, shown in Panel (c). In the NBI phase, the pinch number is largest in the inner core and decreases beyond mid-radius. This behavior closely resembles the profile of the logarithmic density gradient shown in Fig. 4(a) and is consistent with the theoretical concept of the Coriolis momentum pinch, which is found in theoretical [Peeters2007_PRL] and experimental studies [Tala_2009, Tala_2011, Zimmermann_2024] to scale most strongly with the logarithmic density gradient. A similar trend is observed in the NBI+ECRH phase, which exhibits slightly higher $R/L_{n_{e}}$ values in the outer core, reflected in a correspondingly larger pinch number. It should be noted that while the origin of the inward particle pinch emerging during the transition is, first of all, independent of the momentum convection, the modified density gradients obviously affect the Coriolis momentum pinch. In addition, the presented methodology does not disentangle advected and convected momentum. While the observed trends are in line with expectations, the relatively large error bars, together with the possible entanglement of advection and convection, warrant caution in the quantitative interpretation of the obtained results for the pinch number.
@@
The observed modification of the intrinsic torque during the NBI+ECRH phase is consistent with the theoretical concept of residual stress generation arising from profile shearing effects and turbulent eddy tilting during the transition from ITG- to TEM-dominated turbulence [Camenen2011]. In particular, the substantial variations in $R/L_{n_{e}}$, see Fig. 4(a), and in $R/L_{T_{e}}$, see Panel (e), have been identified in previous studies [Angioni_PRL_2011, Grierson_PRL_2017, Hornsby2018, Zimmermann_2024] as key drivers of counter-current intrinsic torque generation. While momentum transport and rotation physics differ fundamentally in stellarators due to the absence of axisymmetry (see discussion in Refs. [hinton1976theory, pp. 289] and [helander2002collisional, pp. 249]), similarities can be identified with observations in the LHD device [Ida_2021_LHD]. In LHD, TEM-dominated scenarios, associated with relatively flat density profiles, exhibit co-current rotation, whereas ITG-dominated scenarios, characterized by density profile peaking in the outer core, tend to exhibit counter-current rotation. This behavior is consistent with the correlation between intrinsic torque and kinetic profile gradients discussed in the present work and in the referenced studies. While this study is focused on H-modes, as L-modes cannot be investigated with the beam modulation technique due to the number of required beams, previous studies on these effects on AUG encompassed both L- and H-modes [Angioni_PRL_2011, McDermott_2014, mcdermott2011effect] and showed that both confinement regimes exhibit the same physical mechanisms of intrinsic rotation generation, hollowing, and sign reversal through the relevant density gradients. In addition, this could also relate to previous work on LHD [Ida2013ReversalTorque], showing changes in the intrinsic torque during the formation of internal transport barriers, which, again, could be correlated with changes in the density gradients.
As in these experiments no scans in the ECRH power were conducted, and as the counter-current intrinsic torque is self-consistently set by the turbulence, mediated by the kinetic profiles, no global threshold for the applied ECRH power is expected to exist for the hollowing. It should be noted that, from a theoretical standpoint [Camenen2011, Hornsby2018], intrinsic torque generation is expected to correlate with second-order derivatives of the kinetic profiles, which are, however, difficult to measure accurately in the experiment. Therefore, first-order derivatives are usually used as a proxy in such an analysis. A definitive interpretation of these observations would require non-linear, global gyrokinetic simulations [Hornsby2018, Wang_PRL_2009, Wang_PoP_2010, Grierson_PRL_2017], which are beyond the scope of the present work. It should be noted that other potential sources of intrinsic torque, such as fast-ion losses, neutral drag, preferential ion losses [Stoltzfus-Dueck_PRL_2012], or neoclassical toroidal viscosity [Zhu_2006, Garofalo_2008, COLE_2011, Shaing_NF_2015], are expected to be negligible in the plasma core for the discharges considered here.
@@
To gain further insight into the underlying turbulence regime, the flux-tube version of the gyrokinetic GKW code [Peeters2004, Peeters2009] was employed using experimental input profiles and an explicit time-integration (initial-value) approach. This allows the identification of the micro-instability with the highest linear growth rate $\gamma$, its dominant mode frequency $\omega_{r}$, and enables an analysis of the mode spectra. These calculations were done at mid-radius for both discharge phases, where the impact of the counter-current intrinsic torque appears to be strongest. The resulting mode spectra are shown in Fig. 8 for the NBI phase in Panel (a) and for the NBI+ECRH phase in Panel (b). In the NBI phase, the mode frequency, see orange points, remains positive across the entire spectrum, indicating turbulence dominated by ITG modes. The growth rates, see blue crosses, are non-zero for smaller $k_{y}\,\rho_{i}$, remain finite over low to intermediate wavenumbers, and decrease to zero towards higher $k_{y}\,\rho_{i}$ (i.e. shorter wavelengths). In contrast, during the NBI+ECRH phase, the spectrum transitions from ITG to TEM (see the sign change in $\omega_{r}$), suggesting the presence of an ITG-TEM mixed-mode regime. This transition is accompanied by reduced $\omega_{r}$ at low $k_{y}\,\rho_{i}$ and increasingly negative values at higher $k_{y}\,\rho_{i}$. The growth rates extend to higher $k_{y}\,\rho_{i}$ and diverge approximately as $\gamma\sim k_{y}^{2}$. However, this increase at high $k_{y}$ is a known feature of such linear calculations in this spectral domain, with higher $k_{y}\,\rho_{s}$ contributing less to the transport. In direct comparison, the growth rates at low $k_{y}\rho_{i}$ are comparable in both phases, while they become larger in the NBI+ECRH phase at higher $k_{y}\rho_{i}$. The absolute values of $\omega_{r}$ are larger in the NBI phase at low $k_{y}\rho_{i}$, whereas at higher $k_{y}\rho_{i}$ they are of similar magnitude and eventually larger in the NBI+ECRH phase.
In summary, and together with the experimental observation of an enhanced inward particle pinch during the NBI+ECRH phase, this provides strong support for the assumption of mixed-mode turbulence, as established by previous works [Angioni_2005, Fable_2008, Fable_2010, Angioni_PRL_2011]. Unfortunately, no turbulence measurements were available during these discharges to further support this modeling. Further gyrokinetic or gyrofluid modeling is left for future work. However, it should be noted that identifying the fastest growing linear modes (e.g. ITG/TEM) remains a meaningful and widely used approach to determine the dominant underlying turbulence regime. In the present case, the experimentally observed core rotation levels are sufficiently small that their impact on the equilibrium $E\times B$ shear, and consequently on mode characteristics, is limited. This supports the consistency of using linear GKW results to identify the instability responsible for the observed counter-current torque.
[…]
Figure 9: Cost function values of the modeling optimization for the NBI+ECRH phase (#29216, 5.57.05.5-7.0 s). The tested pinch number (at mid-radius) is shown on the x-axis and the scaling parameter of the residual stress flux on the y-axis (at mid-radius). The final model solution is indicated by a black marker. The “valley” of acceptable solutions extends predominantly toward positive pinch numbers, suggesting inward convective momentum transport and peaking of the rotation profiles. Positive residual stress values yield a counter-current intrinsic torque contribution.
@@
[…]
Figure 10: Comparison of the momentum fluxes for both discharge phases for #29216 for the discharge phase with dominant NBI heating (2.04.52.0-4.5 s, l.h.s.) and for the phase with NBI+ECRH (5.57.05.5-7.0 s, r.h.s.). The black markers indicate the sum of all momentum flux contributions. Values close to zero demonstrate the temporal stability and self-consistency of the modeled solutions. Positive flux values correspond to an outward flux, which yields an apparent counter-current torque.